# Fourier Circuits in Neural Networks: Unlocking the Potential of Large Language Models in Mathematical Reasoning and Modular Arithmetic

### v1

@@ D.2 Constructions for $\theta^{*}$
  • We define $*\Omega_{q}^{{}^{\prime}*}$ in Definition C.2.
@@
  • The maximum $L_{2,4}$-margin solution $\theta^{*}$ will consist of $16(p-1)$ neurons $*\theta^{*}_{i}\in\Omega_{q}^{{}^{\prime}*}$ to simulate $\frac{p-1}{2}$ type of cosine computation, each cosine computation is uniquely determined a $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$. In particular, for each $\zeta$ the cosine computation is $\cos_{\zeta}(a_{1}+a_{2}+a_{3}-c),\forall a_{1},a_{2},a_{3},c\in\mathbb{Z}_{p}$.
@@
Referencing Lemma C.4, we can identify elements within $\Omega_{q}^{{}^{\prime}}$. Our set $\theta^{*}$ will be composed of $16(p-1)$ neurons, including $32$ neurons dedicated to each frequency in the range $1,\ldots,\frac{p-1}{2}$. Focusing on a specific frequency $\zeta$, for the sake of simplicity, let us use $\cos_{\zeta}(x)$ to represent $\cos(2\pi\zeta x/p)$ and $\sin_{\zeta}(x)$ likewise. We note:
@@
Similarly, all other six terms in Eq. (D.2) can be composed by four neurons with different $(\theta_{u_{1}}^{*},\theta_{u_{2}}^{*},\theta_{u_{3}}^{*},\theta_{w}^{*})$.
@@
The scaling factor $\beta$ for each neuron can be selected such that the entire network maintains an $L_{2,4}$-norm of 1. In this setup, every data point lies exactly on the margin, meaning $q=\operatorname{unif}(\mathbb{Z}_{p})$ uniformly covers points on the margin, thus meeting the criteria for $q^{*}$ as outlined in Definition 3.7. Furthermore, for any input $(a_{1},a_{2},a_{3})$, the function $f$ yields an identical result across all incorrect labels $c^{\prime}$, adhering to Condition 3.10. ∎
D.3 Constructions for $\theta^{*}$ for General $k$ Version
Lemma D.3 (Formal version of Lemma 4.3).
@@ D.3 Constructions for $\theta^{*}$ for General $k$ Version
  • We define $*\Omega_{q}^{{}^{\prime}*}$ in Definition C.2.
@@
  • The maximum $L_{2,k+1}$-margin solution $\theta^{*}$ will consist of $2^{2k-1}\cdot{p-1\over 2}$ neurons $*\theta^{*}_{i}\in\Omega_{q}^{{}^{\prime}*}$ to simulate $\frac{p-1}{2}$ type of cosine computation, each cosine computation is uniquely determined a $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$. In particular, for each $\zeta$ the cosine computation is $\cos_{\zeta}(a_{1}+\dots+a_{k}-c),\forall a_{1},\dots,a_{k},c\in\mathbb{Z}_{p}$.
@@
By Lemma C.4, we can get elements of $*\Omega_{q}^{{}^{\prime}*}$. Our set $\theta^{*}$ will be composed of $2^{2k-1}\cdot{p-1\over 2}$ neurons, including $2^{2k-1}$ neurons dedicated to each frequency in the range $1,\ldots,\frac{p-1}{2}$. Focusing on a specific frequency $\zeta$, for the sake of simplicity, let us use $\cos_{\zeta}(x)$ to represent $\cos(2\pi\zeta x/p)$ and $\sin_{\zeta}(x)$ likewise.
@@
The scaling parameter $\beta$ for each neuron can be adjusted to ensure that the network possesses an $L_{2,k+1}$-norm of 1. For this network, all data points are positioned on the margin, which implies that $q=\operatorname{unif}(\mathbb{Z}_{p})$ naturally supports points along the margin, aligning with the requirements for $q^{*}$ presented in Definition 3.7. Additionally, for every input $(a_{1},\dots,a_{k})$, the function $f$ assigns the same outcome to all incorrect labels $c^{\prime}$, thereby fulfilling Condition 3.10. ∎
Appendix E Check Fourier Frequencies
Section E.1 proves all frequencies are used. Section E.2 proves all frequencies are used for general $k$ version.
E.1 All Frequencies are Used
@@
In this proof, let $j_{1},j_{2},j_{3},j_{4}\in\mathbb{Z}$, and $\theta_{u}=\theta_{u}^{*}\cdot\frac{p}{2\pi}$ to simplify the notation. By Lemma C.4,
@@
$f(a_{1},a_{2},a_{3},c)$ $=\sum_{h=1}^{H}\phi_{h}(a_{1},a_{2},a_{3},c)$
$=\sum_{h=1}^{H}(u_{h,1}(a_{1})+u_{h,2}(a_{2})+u_{h,3}(a_{3}))^{3}w_{h}(c)$
$=(\frac{1}{2p})^{2}\sum_{h=1}^{H}(\cos_{p}(\theta_{u_{h,1}}+\zeta_{h}a_{1})+\cos_{p}(\theta_{u_{h,2}}+\zeta_{h}a_{2})+\cos_{p}(\theta_{u_{h,3}}+\zeta_{h}a_{3}))^{3}\cos_{p}(\theta_{w_{h}}+\zeta_{h}c)$
@@
Setting aside the previously discussed points, it’s established in Lemma B.2 that the function $f$ maintains a consistent margin for various inputs as well as over different classes, i.e., $f$ can be broken down as
@@
E.2 All Frequencies are Used for General $k$ Version
@@
For this proof, for all $j_{1},\dots,j_{k+1}\in\mathbb{Z}$, to simplify the notation, let $\theta_{u}=\theta_{u}^{*}\cdot\frac{p}{2\pi}$, by Lemma C.8, so
@@
Setting aside the previously discussed points, it’s established in Lemma B.2 that the function $f$ maintains a consistent margin for various inputs as well as over different classes, i.e., $f$ can be broken down as
@@
Appendix F Main Result
Section F.1 proves the main result for $k=3$. Section F.2 proves the general $k$ version of our main result.
F.1 Main result for $k=3$
Theorem F.1.
@@ F.1 Main result for $k=3$
By Lemma C.4, we get the single neuron class-weighted margin solution set $*\Omega_{q}^{{}^{\prime}*}$ satisfying Condition 3.10 and $\gamma^{*}$.
By Lemma D.2 and Lemma B.1, we can construct network $\theta^{*}$ which uses neurons in $*\Omega_{q}^{{}^{\prime}*}$ and satisfies Condition 3.10 and Definition 3.7 with respect to $q=\mathrm{unif}(\mathbb{Z}_{p})$. By Lemma B.2, we know it is the maximum-margin solution.
By Lemma E.1, when $j_{1}=j_{2}=j_{3}=-j_{4}\neq 0$, we must have $\widehat{f}(j_{1},j_{2},j_{3},j_{4})>0$. However, as discrete Fourier transform $\widehat{\phi}$ of each neuron is nonzero, for each frequency, we must have that there exists one neuron using it.
F.2 Main Result for General $k$ Version
Theorem F.2 (Formal version of Theorem 4.1).
@@ F.2 Main Result for General $k$ Version
Follow the same proof sketch as Theorem F.1 by Lemma C.8, Condition 3.10, Lemma D.3, Lemma B.1, Definition 3.7, Lemma B.2, Lemma E.2.
Appendix G More Experiments
G.1 One-hidden Layer Neural Network
@@
G.2 One-layer Transformer
In Figure 9 , we train a one-layer transformer with $m=160$ heads attention, i.e., Eq. (10), on $k=3$-sum mod-$p=61$ addition dataset, i.e., Eq. (1). In Figure 10 , we train a one-layer transformer with $m=160$ heads attention, i.e., Eq. (10), on $k=5$-sum mod-$p=17$ addition dataset, i.e., Eq. (1).
Figure 9 and Figure 10 show that the one-layer transformer trained with stochastic gradient descent learns 2-dim cosine shape attention matrices, which is similar to one-hidden layer neural networks in Figure 5 and Figure 7. It means that the attention layer has a similar learning mechanism as neural networks in the modular arithmetic task, where it prefers to learn Fourier-based circuits when trained by SGD.

### v2

@@ F.2 Constructions for $\theta^{*}$
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition E.2.
@@
  • The maximum $L_{2,4}$-margin solution $\theta^{*}$ will consist of $16(p-1)$ neurons $\theta^{*}_{i}\in\Omega_{q}^{{}^{\prime}*}$ to simulate $\frac{p-1}{2}$ type of cosine computation, each cosine computation is uniquely determined a $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$. In particular, for each $\zeta$ the cosine computation is $\cos_{\zeta}(a_{1}+a_{2}+a_{3}-c),\forall a_{1},a_{2},a_{3},c\in\mathbb{Z}_{p}$.
@@
Referencing Lemma E.4, we can identify elements within $\Omega_{q}^{{}^{\prime}}$. Our set $\theta^{*}$ will be composed of $16(p-1)$ neurons, including $32$ neurons dedicated to each frequency in the range $1,\ldots,\frac{p-1}{2}$. Focusing on a specific frequency $\zeta$, for the sake of simplicity, let us use $\cos_{\zeta}(x)$ to represent $\cos(2\pi\zeta x/p)$ and $\sin_{\zeta}(x)$ likewise. We note:
@@
Similarly, all other six terms in Eq. (F.2) can be composed by four neurons with different $(\theta_{u_{1}}^{*},\theta_{u_{2}}^{*},\theta_{u_{3}}^{*},\theta_{w}^{*})$.
@@
The scaling factor $\beta$ for each neuron can be selected such that the entire network maintains an $L_{2,4}$-norm of 1. In this setup, every data point lies exactly on the margin, meaning $q=\operatorname{unif}(\mathbb{Z}_{p})$ uniformly covers points on the margin, thus meeting the criteria for $q^{*}$ as outlined in Definition 3.5. Furthermore, for any input $(a_{1},a_{2},a_{3})$, the function $f$ yields an identical result across all incorrect labels $c^{\prime}$, adhering to Condition 3.8. ∎
F.3 Constructions for $\theta^{*}$ for General $k$ Version
Lemma F.3 (Formal version of Lemma 4.3).
@@ F.3 Constructions for $\theta^{*}$ for General $k$ Version
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition E.2.
@@
  • The maximum $L_{2,k+1}$-margin solution $\theta^{*}$ will consist of $2^{2k-1}\cdot{p-1\over 2}$ neurons $\theta^{*}_{i}\in\Omega_{q}^{{}^{\prime}*}$ to simulate $\frac{p-1}{2}$ type of cosine computation, each cosine computation is uniquely determined a $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$. In particular, for each $\zeta$ the cosine computation is $\cos_{\zeta}(a_{1}+\dots+a_{k}-c),\forall a_{1},\dots,a_{k},c\in\mathbb{Z}_{p}$.
@@
By Lemma E.4, we can get elements of $\Omega_{q}^{{}^{\prime}*}$. Our set $\theta^{*}$ will be composed of $2^{2k-1}\cdot{p-1\over 2}$ neurons, including $2^{2k-1}$ neurons dedicated to each frequency in the range $1,\ldots,\frac{p-1}{2}$. Focusing on a specific frequency $\zeta$, for the sake of simplicity, let us use $\cos_{\zeta}(x)$ to represent $\cos(2\pi\zeta x/p)$ and $\sin_{\zeta}(x)$ likewise.
@@
The scaling parameter $\beta$ for each neuron can be adjusted to ensure that the network possesses an $L_{2,k+1}$-norm of 1. For this network, all data points are positioned on the margin, which implies that $q=\operatorname{unif}(\mathbb{Z}_{p})$ naturally supports points along the margin, aligning with the requirements for $q^{*}$ presented in Definition 3.5. Additionally, for every input $(a_{1},\dots,a_{k})$, the function $f$ assigns the same outcome to all incorrect labels $c^{\prime}$, thereby fulfilling Condition 3.8. ∎
Appendix G Check Fourier Frequencies
Section G.1 proves all frequencies are used. Section G.2 proves all frequencies are used for general $k$ version.
G.1 All Frequencies are Used
@@
In this proof, let $j_{1},j_{2},j_{3},j_{4}\in\mathbb{Z}$, and $\theta_{u}=\theta_{u}^{*}\cdot\frac{p}{2\pi}$ to simplify the notation. By Lemma E.4,
@@
$f(a_{1},a_{2},a_{3},c)$
$=$ $~{}\sum_{h=1}^{H}\phi_{h}(a_{1},a_{2},a_{3},c)$
$=$ $~{}\sum_{h=1}^{H}(u_{h,1}(a_{1})+u_{h,2}(a_{2})+u_{h,3}(a_{3}))^{3}w_{h}(c)$
$=$ $~{}(\frac{1}{2p})^{2}\sum_{h=1}^{H}(\cos_{p}(\theta_{u_{h,1}}+\zeta_{h}a_{1})+\cos_{p}(\theta_{u_{h,2}}+\zeta_{h}a_{2})+\cos_{p}(\theta_{u_{h,3}}+\zeta_{h}a_{3}))^{3}\cos_{p}(\theta_{w_{h}}+\zeta_{h}c)$
@@
Setting aside the previously discussed points, it’s established in Lemma D.2 that the function $f$ maintains a consistent margin for various inputs as well as over different classes, i.e., $f$ can be broken down as
@@
G.2 All Frequencies are Used for General $k$ Version
@@
For this proof, for all $j_{1},\dots,j_{k+1}\in\mathbb{Z}$, to simplify the notation, let $\theta_{u}=\theta_{u}^{*}\cdot\frac{p}{2\pi}$, by Lemma E.8, so
@@
Setting aside the previously discussed points, it’s established in Lemma D.2 that the function $f$ maintains a consistent margin for various inputs as well as over different classes, i.e., $f$ can be broken down as
@@
Appendix H Main Result
Section H.1 proves the main result for $k=3$. Section H.2 proves the general $k$ version of our main result.
H.1 Main result for $k=3$
Theorem H.1.
@@ H.1 Main result for $k=3$
By Lemma E.4, we get the single neuron class-weighted margin solution set $\Omega_{q}^{{}^{\prime}*}$ satisfying Condition 3.8 and $\gamma^{*}$.
By Lemma F.2 and Lemma D.1, we can construct network $\theta^{*}$ which uses neurons in $\Omega_{q}^{{}^{\prime}*}$ and satisfies Condition 3.8 and Definition 3.5 with respect to $q=\mathrm{unif}(\mathbb{Z}_{p})$. By Lemma D.2, we know it is the maximum-margin solution.
By Lemma G.1, when $j_{1}=j_{2}=j_{3}=-j_{4}\neq 0$, we must have $\widehat{f}(j_{1},j_{2},j_{3},j_{4})>0$. However, as discrete Fourier transform $\widehat{\phi}$ of each neuron is nonzero, for each frequency, we must have that there exists one neuron using it.
H.2 Main Result for General $k$ Version
Theorem H.2 (Formal version of Theorem 4.1).
@@ H.2 Main Result for General $k$ Version
Follow the same proof sketch as Theorem H.1 by Lemma E.8, Condition 3.8, Lemma F.3, Lemma D.1, Definition 3.5, Lemma D.2, Lemma G.2.
Appendix I More Empirical Details and Results
I.1 Implement Details
Licenses for Existing Assets & Open Access to Data and Code. Our code is based on a brilliant open source repository, https://github.com/Sea-Snell/grokking, which requires MIT License. We provide all of our codes in the supplemental material, including dataset generation code. We do not require open data access as we run experiments on synthetic datasets, i.e., modular addition.
Experimental Result Reproducibility. We provide all of our codes in the supplemental material with a clear README file and clear configuration files for our experiments reproducibility.
Experimental Setting/Details & Experiment Statistical Significance. The detailed configuration can be found in supplemental material. We make a copy version here for convenience.
For two-layer neural network training, we have the following details:
  • number of data loader workers: $4$
  • batch size: $1024$
  • learning rate: $5\times 10^{-3}$
  • regularization strength $\lambda$: $0.005$
  • AdamW hyper-parameter $(\beta_{1},\beta_{2})$: $(0.9,0.98)$
  • warm-up steps: $10$
For one-layer Transformer training, we have the following details:
  • number of data loader workers: $4$
  • batch size: $1024$
  • learning rate: $1\times 10^{-3}$
  • regularization strength $\lambda$: $0.001$
  • AdamW hyper-parameter $(\beta_{1},\beta_{2})$: $(0.9,0.98)$
  • warm-up steps: $10$
All results we ran 3 times with different random seeds. In Figure 4, we reported the mean and variance range.
Experiments Compute Resources. All experiments is conducted on single A100 40G NVIDIA GPU. All experiments can be finished in at most three days.
I.2 One-hidden Layer Neural Network
@@
I.3 One-layer Transformer
In Figure 9 , we train a one-layer transformer with $m=160$ heads attention, on $k=3$-sum mod-$p=61$ addition dataset, i.e., Eq. (1). In Figure 10 , we train a one-layer transformer with $m=160$ heads attention, on $k=5$-sum mod-$p=17$ addition dataset, i.e., Eq. (1).
Figure 9 and Figure 10 show that the one-layer transformer trained with stochastic gradient descent learns 2-dim cosine shape attention matrices, which is similar to one-hidden layer neural networks in Figure 5 and Figure 7. This means that the attention layer has a similar learning mechanism to neural networks in the modular arithmetic task, where it prefers to learn Fourier-based circuits when trained by SGD.